Item7 Item 7 Your firm is a Swiss importer of bicycles. You have placed an order...

Item7

Item 7

Your firm is a Swiss importer of bicycles. You have placed an order with an Italian firm for €1,000,000 worth of bicycles. Payment (in euro) is due in 12 months. Use a money market hedge to redenominate this one-year receivable into a Swiss franc-denominated receivable with a one-year maturity.

Contract Size Country U.S. $ equiv. Currency per U.S. $
£ 10,000 Britain (pound) $ 1.9600 £ 0.5102 interest APR
12 months forward $ 2.0000 £ 0.5000 rates
10,000 Euro $ 1.5600 0.6410 i$ = 1 %
12 months forward $ 1.6000 0.6250 i = 2 %
SFr. 10,000 Swiss franc $ 0.9200 SFr. 1.0870 i£ = 3 %
12 months forward $ 1.0000 SFr. 1.0000 iSFr. = 4 %

The following were computed without rounding. Select the answer closest to yours.

Multiple Choice

  • SFr. 1,544,705.88

  • SFr. 1,600,000

  • SFr. 800,000

  • SFr. 1,728,900.26
Answer #1

Step 1: Borrow Euro 1 Million at APR of 2%

The Present Value (PV) of Euro 1 Million at APR of % = Future value / (1+Rate of Return) ^n

n= No. of Years.= 12 months or 1 year (as given in question)

Future value = Euro 1 Million of Euro 1,000,000

"r" of Rate of Return (APR or Euro ) = 2%

Present Value (PV) = Euro 1,000,000/(1+2%)^1

=> 1,000,000/ (1.02)

Therefore, Present Value of Euro 1,000,000 = Euro 980,392.16

Therefore, if we borrow Euro 1,000,000 at APR of 2 %, we will received Euro 980,392.16

Step 2: On the date of maturity of borrowing, we will use the receivable of Euro 980,392.16 (see Step 1) to repay the Loan.

Step 3: We will then sell Euro 980,392.16 and receive in US $ at the current exchange rate of $ 1.56:

Euro 980,392.16 * $ 1.56/ Euro 1

=> Euro 980,392.16 * 1.56

=> US $ 15,29,411.76

Therefore, amount of US $ received on sale = $ 1,529,411.76

Step 4: Buy Swiss Franc by selling $ 1,529,411.76 at Exchange rate of US $0.92/ 1 CHF (CHF = Swiss Franc)

We will receive, 15,29,411.76*1/0.92 ( if 1 CHF = $0.92; then US $ 1,529,411.76) = CHF 1,662,404.09.

So, we will need to buy CHF 1,662,404.09

Step 5: Re-denominate the CHF of 1,662,404.09 into a 1 year Receivable at APR of 4 % (given in question)

So, the 1 year Receivable will be equal to CHF 1,662,404.09 * (1+r%)^1

r = APR of 4%

-=> CHF 1,662,404.09 * 1.04 = CHF 1,728,900.26

Answer : The One year CHF Receivable will be = CHF 1,728,900.26